OpenStudy (anonymous):

1 over the cube root of the quantity of x to the negative sixth power

OpenStudy (yanasidlinskiy):

@Basher that is unnecessary. @grray7 \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)

OpenStudy (anonymous):

good reply basher

OpenStudy (bibby):

recall that fractional exponents are roots like \(a^{1/2}=\sqrt{a}\) \(\large a^{1/4}=\sqrt[4]{a}\) and that powers raised to powers get multiplied like \((a^m)^n=a^{m*n}\)

OpenStudy (yanasidlinskiy):

\(\large \sqrt[3]{x^6} = x^2\)

OpenStudy (yanasidlinskiy):

Understand?

OpenStudy (anonymous):

yeah, thanks. And BTW's I like your profile. :)

OpenStudy (yanasidlinskiy):

Your Welcome!!! Mine?

OpenStudy (anonymous):

yes, @YanaSidlinskiy

OpenStudy (bibby):

\(\huge \frac{1}{\sqrt[3]{x}^{-6}}=\sqrt[3]{x}^{6}=x^{\frac{1}{3}^6}\)

OpenStudy (anonymous):

Thanks Bibby

OpenStudy (yanasidlinskiy):

Lol, thanx!! @grray7

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

@yanasidlinskiy

OpenStudy (yanasidlinskiy):

Did u need me? grray7

OpenStudy (anonymous):

I do now, if you could help me out with \[\sqrt{-3x-10}=x\]